# CRYSP: a Total-Body PET based on cryogenic cesium iodide crystals

### v1

@@ VII-B Spatial resolution
512 (1/2 of FOV) 10 1.68 1.71 1.71 3.05 3.12 3.12
128 (1/8 of FOV) 10 1.68 1.71 1.71 3.06 3.11 3.11
512 (1/2 of FOV) 100 1.70 1.72 1.76 3.09 3.12 3.21
128 (1/8 of FOV) 100 1.72 1.71 1.76 3.12 3.12 3.21
512 (1/2 of FOV) 200 1.77 1.74 1.90 3.22 3.17 3.45
128 (1/8 of FOV) 200 1.85 1.80 1.99 3.36 3.28 3.62
VII-C Count rates and NECR
The count rates performance of the device is calculated by simulating a solid circular cylinder composed of polyethylene with an outside diameter of 203 mm, an overall length of 700 mm, and an internal volume of 22,000 mL. The cylinder is placed in the center of the FOV and axially aligned with the scanner. A 6.4 mm hole is added parallel to the central axis of the cylinder, at a radial distance of 4 mm, according to the NEMA specifications. Inside the hole, a tubing of length 700 mm, inside diameter of 3.2 mm, and outside diameter of 4.8 mm was inserted. In the simulation, $4.5\times 10^{6}$ positrons were generated uniformly within the tubing volume. The time window was varied in order to simulate different activities, ranging from 0 to 40 kBq/mL.
Rates of total, true, scatter, and noise-equivalent counts were calculated as detailed in section 4 of the NEMA NU 2-2018 protocol and are shown in fig. 16. A peak NECR of 1.79 Mcps is obtained at an equivalent concentration of 14 kBq/mL.
Figure 16: True (black solid line), scatter (black dashed line), total (black dotted line), NECR (red solid line) and random (blue solid line) rates as a function of activity in the phantom tubing.
VII-D Pile-up
@@
In order to quantify the effect of pile-up more realistically, we simulated a simplified MIRD human phantom [41] inside CRYSP1M, using approximate relative activities in each organ as detailed in ref.[42]. Two scenarios are being considered: one where the patient’s entire upper body is inside the scanner and another where the head is outside the FOV. In our simulation, signals pass through the pile-up processor described in section IV-B. An event is classified as pile-up if the processor fails to resolve a second signal arriving within the integration window of the first. Fig. 17 illustrates the fraction of events with unresolved pile-up as a function of the activity administered to the patient. For a typical PET administered activity of 300 MBq, approximately 20% of events fall into this category. This fraction can be halved by positioning the patient’s head, where a significant portion of the activity is concentrated, outside the FOV.
@@
A key advantage of TBPETs, however, is their ability to maintain image quality with up to an order of magnitude lower administered dose, as demonstrated by uEXPLORER [43]. At an activity level of 30 MBq, this results in an almost negligible fraction of unresolved pile-up, approximately 2%.
@@
VIII Image reconstruction in CRYSP
@@
The quality of the images reconstructed by CRYSP represents the chief parameter to be compared with existing solutions, since it encodes sensitivity, spatial resolution, time resolution, energy resolution and NECR performances. To do so, we compare the contrast recovery coefficient (CRC) of CRYSP with the one of a simulated pixelated LYSO PET of the same size (774 mm diameter, 1024 mm length and 2 $X_{0}$ as thickness). The pixels are assumed to have a transverse size of $3\times 3$ mm2. The CTR of the simulated LYSO PET is set to 350 ps, halfway between the performances of Quadra [4] and uEXPLORER [43], which enables TOF reconstruction.
@@
In order to better assess the impact of Compton scattering with the patient’s body, we simulated a modified Jaszczak phantom with an inner diameter of 266 mm (instead of the typical 216 mm) and an axial length of 686 mm (instead of 186 mm), placed at the center of the AFOV as shown in fig. 12. Adopting a phantom with larger length and diameter allows for a more accurate estimation of Compton scattering effects in human patients. This is particularly important in a FBPET, where a significant fraction of detected gammas is emitted at low angles and passes through substantial portions of the body, increasing the likelihood of Compton scattering. The simulation was performed with $3\times 10^{9}$ electron-positron annihilation events distributed in 10 s of time, corresponding to an activity concentration of 7.9 kBq/mL. This value, although smaller than the peak NECR, has been chosen because it gives a manageable pile-up fraction of less than 20%.
The CRC was calculated for all rod radii, both for rods placed near the center of the phantom (inner rods) and for rods placed closer to the surface (outer rods), as shown in fig. 18. Reconstruction has been performed with the parallelproj Python package [44], which allows to run a highly-parallelized implementation of the MLEM algorithm on GPUs. The reconstruction voxel size was set to 1 mm in all three axes, a value close to the spatial resolution obtained by the CNN, and the algorithm was run for 10 iterations. Normalization and attenuation corrections were applied in both cases and no point-spread function modeling was used.
The plots show that a PET scanner based on monolithic CsI crystals has performances that can be compared with one employing LYSO pixels. Most interestingly, the CRC for rods placed near the center of the phantom is significantly better with CsI: in this case, in fact, the probability for a gamma to undergo Compton scattering in the phantom material is larger, and the extremely good energy resolution of cryogenic CsI more than compensates for the lower sensitivity and lack of TOF. For the inner rods, the CRC goes from 21% for the smallest radius (1.60 mm) to 96% for the largest one (5.55 mm). In the case of the outer rods, the variation is smaller, from 39% to 68%.
@@
Figure 18: Top: reconstructed images with a modified Jaszczak phantom for a PET scanner with monolithic cryogenic CsI crystals (left) and pixelated LYSO crystals (right). Normalization and attenuation corrections are applied and the color scale is the same. Bottom: contrast recovery coefficients as a function of the radius of the inner (solid) and outer (dashed) rods. The regions being considered corresponds to the red circumferences shown above.
IX Summary and future prospects
@@
This work introduces CRYSP, a novel PET system based on cryogenic pure CsI monolithic crystals with the potential to achieve high performances at reduced costs. By operating at cryogenic temperatures, CsI offers a light yield of approximately $10^{5}$ photons/MeV, which translates into an energy resolution below 7% FWHM. Despite its relatively slow scintillation decay time, this material can achieve a coincidence time resolution of approximately 1.5 ns, sufficient for PET applications.
@@
Simulation studies show the ability to achieve millimeter-scale spatial resolution for interaction vertex reconstruction within a monolithic cryogenic CsI crystal. Overall, the simulated CRYSP1M scanner achieves sensitivity, NECR, and CRCs comparable to existing solutions of the same size, despite the absence of TOF capabilities. Notably, CRYSP1M performs better than a pixelated LYSO PET scanner in situations where energy resolution is critical, such as in samples with a large number of Compton-scattered gammas.
@@
The use of cryogenics, although standard in medical diagnosis (e.g., nuclear magnetic resonance, NMR) is new in PET devices. However, and unlike NMR, CRYSP cryogenics only requires a bath of N2, an thus conventional, low-cost techniques for cooling and signal extraction (e.g., cheap feedthroughs) can be deployed. As a result, the cryogenics increases the total cost of the device by less than 5% and can be implemented (again, like in the case of NMR) transparently to the final user.
In summary, CRYSP may offer a cost-effective and high-performing alternative for TBPET scanners, holding great promise for wider accessibility and utility in both clinical and research settings.
Acknowledgments
@@
SRS acknowledges the support of a fellowship from “la Caixa Foundation” (ID 100010434) with code LCF/BQ/PI22/11910019 and of a grant from the Basque Country Government with code PUE24-10.
JR acknowledges support from the Generalitat Valenciana of Spain under grant CIDEXG/2023/16.

### v2

@@ 3.2 Spatial resolution
CRYSP 512 (1/2 of FOV) 10 3.35 3.42 3.42 6.13 6.26 6.26
128 (1/8 of FOV) 10 3.36 3.41 3.41 6.15 6.24 6.25
LYSO 512 (1/2 of FOV) 10 3.52 3.61 3.61 6.44 6.60 6.60
128 (1/8 of FOV) 10 3.51 3.62 3.61 6.39 6.60 6.58
CRYSP 512 (1/2 of FOV) 100 3.39 3.43 3.52 6.21 6.28 6.44
128 (1/8 of FOV) 100 3.43 3.42 3.52 6.28 6.26 6.44
LYSO 512 (1/2 of FOV) 100 4.21 3.72 3.71 7.68 6.78 6.76
128 (1/8 of FOV) 100 4.26 3.99 3.78 7.77 7.27 6.89
CRYSP 512 (1/2 of FOV) 200 3.53 3.47 3.79 6.46 6.17 6.94
128 (1/8 of FOV) 200 3.69 3.60 3.97 6.76 6.59 7.27
LYSO 512 (1/2 of FOV) 200 4.89 4.61 3.94 8.91 8.40 7.18
128 (1/8 of FOV) 200 5.01 4.54 3.91 9.13 8.27 7.13
3.3 Count rates and NECR
The count‐rate performance was evaluated for both the CRYSP and the reference LYSO‐based PET scanner using the NEMA NU 2–2018 protocol. A solid circular cylinder made of polyethylene with an outer diameter of 203 mm, an overall length of 700 mm, and an internal volume of 22,000 mL was simulated. The phantom was placed at the center of the FOV and axially aligned with the scanner. A 6.4 mm hole, positioned 4 mm off the central axis, contained a 700 mm-long tubing (inner/outer diameters of 3.2/4.8 mm) filled with a uniform positron source. In each configuration, $4.5\times 10^{6}$ positrons were generated uniformly within the tubing volume, and the acquisition time window was varied to reproduce different activity concentrations between 0 and 40 kBq/mL.
Rates of total, true, scattered, and noise-equivalent counts were calculated according to Section 4 of the NEMA NU 2–2018 standard and are shown in Fig. 10. A peak NECR of $(1.79\pm 0.04)$ Mcps was obtained for CRYSP at an equivalent activity concentration of $14.0$ kBq/mL, while the LYSO-based PET reached $(2.97\pm 0.05)$ Mcps at $25.6$ kBq/mL, in good agreement with the reported measurement from Quadra [prenosil2022performance] (2.96 Mcps for MRD at 27.5 kBq/mL). The corresponding scatter fraction (SF) was found to be approximately 20% for CRYSP and 40% for the LYSO-based PET, as shown in Fig. 10 (right), indicating that the improved energy resolution of cryogenic CsI effectively reduces scattered coincidences by about a factor of two. Also in this case, the SF for the LYSO-based simulated PET is in good agreement with the Quadra reported measurement [prenosil2022performance].
Figure 10: Left: true (solid line), scatter (dashed line), random (dash-dotted line), and total (dotted line), and NECR curve (red line) for the CRYSP system. Right: NECR curves (red, left axis) and scatter fractions (black, right axis) for CRYSP (solid), the Biograph Vision Quadra [prenosil2022performance] (dashed), and a LYSO-based simulated PET scanner (dotted). Both plots are as a function of the activity concentration in the phantom tubing.
3.4 Pile-up
@@
In order to quantify the effect of pile-up more realistically, we simulated a simplified MIRD human phantom [parach2011assessment] inside CRYSP1M, using approximate relative activities in each organ as detailed in ref. [Dias2022NormalVF]. Two scenarios are being considered: one where the patient’s entire upper body is inside the scanner and another where the head is outside the FOV. In our simulation, signals pass through the pile-up processor described in section 2.3.2. An event is classified as pile-up if the processor fails to resolve a second signal arriving within the integration window of the first. Fig. 11 illustrates the fraction of events with unresolved pile-up as a function of the activity administered to the patient. For a typical PET administered activity of 300 MBq, approximately 20% of events fall into this category. This fraction can be halved by positioning the patient’s head, where a significant portion of the activity is concentrated, outside the FOV.
@@
A key advantage of TBPETs, however, is their ability to maintain image quality with up to an order of magnitude lower administered dose, as demonstrated by uEXPLORER [badawi2019first]. At an activity level of 30 MBq, this results in an almost negligible fraction of unresolved pile-up, approximately 2%. For comparison, in the reference LYSO-based PET scanner, the much shorter scintillation decay time and smaller detector block area result in a negligible pile-up fraction under the same activity conditions.
@@
3.5 Image reconstruction in CRYSP
@@
The quality of the images reconstructed by CRYSP represents the chief parameter to be compared with existing solutions, since it encodes sensitivity, spatial resolution, time resolution, energy resolution and NECR performances. To do so, we compare the contrast recovery coefficient (CRC) and signal-to-noise ratio (SNR) of CRYSP with the one of the simulated pixelated LYSO PET, as described in section 2.4, with and without TOF capabilities.
@@
In order to better assess the impact of Compton scattering with the patient’s body, we simulated a modified Jaszczak phantom with an inner diameter of 266 mm (instead of the typical 216 mm) and an axial length of 686 mm (instead of 186 mm), placed at the center of the AFOV. Adopting a phantom with larger length and diameter allows for a more accurate estimation of Compton scattering effects in human patients. This is particularly important in a TBPET, where a significant fraction of detected gammas is emitted at low angles and passes through substantial portions of the body, increasing the likelihood of Compton scattering. The simulation was performed with $3\times 10^{9}$ electron-positron annihilation events distributed in 10 s of time, corresponding to an activity concentration of 7.9 kBq/mL. This value, although smaller than the peak NECR, has been chosen because it gives a manageable pile-up fraction of less than 20%.
@@
Figure 12: Top: reconstructed images with a modified Jaszczak phantom for a PET scanner with monolithic cryogenic CsI crystals (left) and pixelated LYSO crystals with TOF (center) and without TOF (right). Normalization and attenuation corrections are applied and the color scale is the same. Bottom: contrast recovery coefficients and signal-to-noise ratios as a function of the radius of the inner (solid) and outer (dashed) rods. The regions being considered correspond to the red circumferences shown above.
The CRC and SNR were calculated for all rod radii, both for rods placed near the center of the phantom (inner rods) and for rods placed closer to the surface (outer rods), as shown in fig. 12. Reconstruction has been performed with the parallelproj Python package [schramm2024parallelproj], which allows to run a highly-parallelized implementation of the MLEM algorithm on GPUs. The reconstruction voxel size was set to 1 mm in all three axes, a value close to the spatial resolution obtained by the CNN, and the algorithm was run for 10 iterations, where the CRC and SNR values plateau. Normalization and attenuation corrections were applied in both cases and no point-spread function modeling was used. A listmode MLEM algorithm was employed for both the CRYSP and the LYSO-based PET reconstructions. All oblique LORs arising from the 100 cm axial coverage were included without any axial compression or sinogram rebinning. No time-of-flight (TOF) information was used in the CRYSP case, given its CTR of 1.5 ns, which is insufficient to provide a meaningful TOF constraint. For the LYSO-based PET system, reconstructions were performed both with TOF information (assuming a 350 ps CTR) and without TOF, in order to assess the impact of TOF on image quality. Voxel size and projector settings were kept identical across the two cases.
The plots show that a PET scanner based on monolithic CsI crystals has performances that can be compared with ones employing LYSO pixels. Most interestingly, the CRC for rods placed near the center of the phantom is significantly better with CsI: in this case, in fact, the probability for a gamma to undergo Compton scattering in the phantom material is larger, and the significantly better energy resolution of cryogenic CsI and its improved d.o.i. resolution more than compensates for the lower sensitivity and lack of TOF. For the inner rods, the CRC goes from 21% for the smallest radius (1.60 mm) to 96% for the largest one (5.55 mm). In the case of the outer rods, the variation is smaller, from 39% to 68%. As expected, the pixelated LYSO-based PET reconstruction without TOF performs worse than its TOF-enabled counterpart. Moreover, due to the combination of poorer energy resolution and the absence of d.o.i. information, its performance is also inferior to that of CRYSP.
4 Discussion
The development of Total-Body PET (TBPET) apparatus represents a major leap forward in medical imaging. On one hand, it enables true multi-modal and multi-organ imaging; on the other, it allows for significant reductions in radiation dose and/or acquisition time. These improvements open new possibilities, such as routine PET imaging for children, and enable faster patient throughput.
While early demonstrators, such as the uEXPLORER scanner [cherry2018total], achieve axial lengths of up to 200 cm, excellent performance can also be obtained with smaller coverage, with 100 cm representing a good compromise for many applications, as demonstrated by the Biograph Vision Quadra [prenosil2022performance].
However, the high costs associated with TBPETs may limit their widespread adoption. These costs are dominated by the high price of LYSO, the material of choice for virtually all modern PET scanners. LYSO is a rare earth material, and both its production and supply are highly concentrated geographically, raising concerns about future availability and cost.
Another challenge is that increasing the axial length of a PET scanner faces a problem of diminishing returns, linked to two main issues: (i) increased multiple scattering as photons travel longer distances through the patient’s body, and (ii) the parallax effect, which becomes a major source of reconstruction error for photons traveling at small angles with respect to the $z$-axis.
The impact of multiple scattering can be mitigated by excellent energy resolution, allowing the selection of photons with energies close to 511 keV, thereby reducing scattering. The parallax effect, in turn, can be significantly suppressed by using monolithic crystals combined with modern reconstruction techniques, such as those based on convolutional neural networks (CNNs). Monolithic crystals also offer a substantial cost reduction compared with their pixelated counterparts.
The balance between d.o.i., TOF, and scatter fraction can be analyzed in terms of their respective contributions to the SNR. The SNR gain provided by TOF information increases with the patient diameter approximately as $\sqrt{D/\Delta x_{\mathrm{TOF}}}$ [budinger1983time, vandenberghe2016recent], where $\Delta x_{\mathrm{TOF}}=c\,\Delta t/2$. For a 350 ps CTR ($\Delta x_{\mathrm{TOF}}\approx 5.25$ cm) and a 20 cm NEMA NECR phantom, this corresponds to an SNR improvement of approximately $\times 1.95$ relative to a non-TOF system.
Information on the d.o.i. in CRYSP reduces parallax broadening, resulting in a narrower and more spatially uniform point spread function. While this does not necessarily translate into a global increase in the image SNR, the better spatial resolution can reduce the partial volume effect [mannheim2023characterization], thus improving the local SNR around small structures by limiting the spread of counts into neighboring voxels. In addition, CRYSP exhibits a substantially lower scatter fraction (approximately 20% versus 40% for the LYSO-based configurations), corresponding to an expected SNR benefit of approximately $\sqrt{(1-0.2)/(1-0.4)}\approx 1.15$. Consistent with these considerations, Fig. 12 shows that CRYSP achieves SNR values comparable to those of the pixelated LYSO-based PET across the examined rods, effectively compensating for the absence of TOF and clearly outperforming the non-TOF LYSO configuration.
These considerations point to the possibility of developing a low-cost TBPET system based on an abundant and inexpensive scintillator: CsI. The key realizations are that: (a) CsI must operate at cryogenic temperatures, where it offers a light yield three times higher than LYSO and an energy resolution roughly twice as good; and (b) the system should use monolithic crystals, which combine lower cost with superior performance at large incidence angles.
The cost of monolithic CsI can be up to an order of magnitude lower than that of pixelated LYSO, while cryogenic operation (in the well-understood regime of liquid nitrogen) adds no more than about 5% to the cost of a conventional PET scanner. Importantly, 3D position reconstruction via neural networks in monolithic crystals provides a complete solution to the parallax problem.
[…]
The results presented here demonstrate that the overall performance of a PET scanner based on monolithic CsI crystals can be competitive with that of much more expensive and complex LYSO-based systems. This offers a viable path to implementing TBPET technology at a cost affordable to most hospitals.
5 Conclusion
@@
This work introduces CRYSP, a novel PET system based on cryogenic pure CsI monolithic crystals, with the potential to achieve high performance at significantly reduced cost. Operating at cryogenic temperatures, CsI provides a light yield of approximately $10^{5}$ photons/MeV, resulting in an energy resolution better than 7% FWHM [Soleti:2024flw]. Despite its relatively slow scintillation decay time, it can achieve a coincidence timing resolution of about 1.5 ns, sufficient for PET applications.
@@
Simulation studies indicate millimeter-scale spatial resolution for interaction vertex reconstruction within a monolithic cryogenic CsI crystal. Overall, the simulated CRYSP1M scanner achieves sensitivity, NECR, and CRCs comparable to existing systems of similar size, despite the absence of TOF capabilities. Notably, CRYSP1M has the potential to outperform a pixelated LYSO PET scanner in situations where energy resolution is critical, such as imaging samples with a high fraction of Compton-scattered gamma rays.
@@
While cryogenics is standard in other medical imaging modalities (e.g., MRI), its application in PET devices is new. In contrast to MRI, CRYSP cryogenics requires only a liquid nitrogen bath, enabling the use of conventional, low-cost cooling and signal-extraction techniques (e.g., inexpensive feedthroughs). As a result, cryogenics adds less than 5% to the overall device cost and can be implemented transparently for the end user.
In summary, CRYSP offers a promising, cost-effective, high-performance alternative for TBPET scanners, with the potential to significantly expand access to advanced PET imaging in both clinical and research contexts.
[…]
@@
[…]
SRS acknowledges the support of a fellowship from “la Caixa Foundation” (ID 100010434) with code LCF/BQ/PI22/11910019 and of a grant from the Basque Country Government with code PUE24-10. JR acknowledges support from the Generalitat Valenciana of Spain under grant CIDEXG/2023/16.
[…]
The data that support the findings of this study are available upon reasonable request from the authors.